OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? y = 4x - 1 y = 4x - 5 y = 4x - 10 y = 4x + 1

OpenStudy (anonymous):

Also please tel me how to solve this. Thanks!

OpenStudy (raden):

the equation of straight line : y = mx + b with m=slope, b=constant

OpenStudy (raden):

u just subtitute the m=4 and and the coordinat point (x,y)=(3,2) to formula above

OpenStudy (anonymous):

I still can't figure it out.

OpenStudy (raden):

y = mx + b first, put m=4, it means y = 4x + b, right ?

OpenStudy (raden_zaikaria):

r u there?

OpenStudy (raden):

then to find the value of b, subtitute the point (3,2) to equation above : y = 4x + b 2 = 4*3 + b 2 = 12 + b or b = 2 - 12 b = -10 so, y = 4x + b y = 4x -10

OpenStudy (anonymous):

I am here.

OpenStudy (anonymous):

So the slope is always M? And in this case M = 4? Correct?

OpenStudy (anonymous):

I get it, you explained it very well. Thanks!